the equation x^2+bx+c=0 has roots -5 and 7. find the value of b and the value of c.


1
Expert's answer
2022-01-05T18:04:56-0500

"x^2+bx+c=(x+5)(x-7)."

"x^2+bx+c=x^2+(5-7)x-5*7."

"x^2+bx+c=x^2-2x-35."

"b=-2, \;c=-35."